Find the fifth term in expression(x+2w)^10
Answer by stanbon(75887)

Find the fifth term in expression(x+2w)^10
-------
That term will have x^6 in it.
-----
5th term: 10C6*x^6*(2w)^4 = 210*x^6(16w^4) = 3360x^6w^4
====================================
Cheers,
Stan H.
====================================